OpenStudy (anonymous):

Write the equation for the locus of points 3 units from the point ( 0,1) .

OpenStudy (anonymous):

i gave you the equation of a circle earlier i think x^2+(y-1)^2=3^2

OpenStudy (anonymous):

so for one unit from the point ( -1,0) it would be (x-1)^2 + (y-0)^2=1 ? @dockworker ?

OpenStudy (anonymous):

(x+1)^2+y^2=1^2

OpenStudy (anonymous):

(h,k) is the center, but there is a minus sign in front of h and k in the equation of a circle, so if h or k is negative, change to addition sign

Directrix (directrix):

Is this an "in the plane" or an "in space" locus problem?

OpenStudy (anonymous):

oh and noooo ....

OpenStudy (anonymous):

thanks again @dockworker

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

in space ? what does that mean ? @directrix

Directrix (directrix):

I was thinking about the problem: What is the locus of points 3 units from a given point? In 2-D, the locus is a circle. In 3-D, it is a sphere. In your problem, the coordinates of the given point are in 2-D (just two coordinates for the point) so I'm thinking that the locus is a circle in 2-D with equation as has been settled upon.

OpenStudy (anonymous):

oh ok ... well my teacher hasnt taught me that yet ''/ @directrix

Directrix (directrix):

Okay. Over at the link given below, scroll down a bit and you can get an idea of the 2-D and 3-D difference in such a problem. http://www.mathsisfun.com/sets/set-of-points.html

OpenStudy (anonymous):

@directrix thanks for the link and for introducing me to something new !!